Each page of a book measures 12 inches by 8 inches. The cover extends beyond the shorter side one half inch in both directions. What is the permeter of the book's cover in inches?.
Answer by fractalier(6550)

If the cover extends beyond the short side by 1/2" on each side, it must be 13" x 8".
Perimeter P = 2L + 2W = 2(13) + 2(8) = 26 + 16 = 42 inches